RC meeting and Run Plan, Thursday February 23th

From SBS wiki
Jump to: navigation, search

RC Report

Summary

  • Beam Delivery until middle of Owl was reasonable
  • Around 4:00am started having problems with the BigBite Magnet
  • Techs recovered BigBite Magnet during morning
  • Working on Beam Recovery

Recent Issues

  • BigBite Magnet trips with overtemperature and overvoltage error, now resolved

Current Status

  • Beam Recovery, then return to production

Run Plan

  • Continue running 45 uA on He3


Data Collection and Polarization Status

  • 39.7 C of uncorrected charge for this kinematic (44.9C of corrected charge which is 28% of the plan)
  • Last NMR 51.5%

Other Topics/Status Updates

  • Readiness target replacement
  • Moller polarimeter readiness
  • 7 Shifts still need filled during the remaining run period

Requested Accesses to the Hall